# Enhanced Ground–Satellite Direct Access via Onboard Rydberg Atomic Quantum Receivers

### v1

@@ Abstract
Ground-satellite links for 6G networks face critical challenges, including severe path loss, tight size‑weight‑power limits, and congested spectrum, all of which significantly hinder the performance of traditional radio frequency (RF) front ends. This article introduces the Rydberg Atomic Quantum Receiver (RAQR) for onboard satellite systems, a millimetre‑scale front end that converts radio fields to optical signals through atomic electromagnetically induced transparency. RAQR’s high sensitivity and high frequency selectivity address link‑budget, payload, and interference challenges while fitting within space constraints. A hybrid atomic–electronic design and supporting signal model demonstrate enhanced data rate, coverage, and sensing accuracy relative to conventional RF receivers. The article concludes with integration strategies, distributed‑satellite concepts, and open research problems for bringing RAQR‑enabled satellite payloads into service.
@@ I Introduction
Spectrum Congestion and Interference Vulnerability: Satellite frequency bands are heavily shared among civil, commercial, and governmental users (e.g., L, S, C, Ku bands). To increase throughput, aggressive frequency reuse strategies are common. These strategies, in turn, introduce substantial co-channel and adjacent-channel interference from other ground stations or satellite systems [5]. Moving to higher frequency bands (e.g., Ka or Q/V bands) offers more bandwidth but brings new design challenges for onboard receivers [6]. These challenges include increased rain fade, stricter hardware linearity requirements, and elevated thermal instability in front-end components. Furthermore, ground-based jamming poses a direct threat to uplink integrity, which requires onboard receivers with robust interference rejection.
Overcoming these intertwined constraints calls for a receive architecture on the satellite that combines very high sensitivity, compact physical dimensions, and minimal power consumption while maintaining interference tolerance. This motivates the investigation of radically different front-end architectures capable of operating under extreme power, size, and spectral isolation constraints, particularly for the reception of ground-originated signals. Quantum-enhanced sensing technologies represent a particularly promising direction. One such approach is the Rydberg atomic quantum receiver (RAQR), which utilizes highly excited Rydberg states of atoms to transduce RF fields from the ground segment into optical signals [7, 8]. These receivers can directly demodulate RF signals via electromagnetically induced transparency (EIT) without requiring conventional components such as local oscillators or RF mixers. Their unique properties, such as wide frequency tunability, ultra-high field sensitivity to weak signals, inherent optical isolation from thermal noise, and the potential for chip-scale integration, make RAQRs an intriguing candidate for next-generation spaceborne front-ends [7].
@@
TABLE I: RAQR Solutions to Key Satellite Communication Challenges
Satellite Communication Challenges RAQR-Based Solutions
Link Budget Constraints: Severe FSPL (150 - 200 dB), constrained power, and limited antenna gain. High Sensitivity: Ultra-high dipole moment between Rydberg transitions and limited thermal noise impacts.
Payload SWaP Constraints: Constrained volume and mass budgets, limited solar panel area and power generation. Miniaturized Architecture: Flexible extension for multi-functional receivers using a single vapor cell, wavelength-independent sensing volume, and no traditional RF frontend.
Spectrum Congestion and Interference: Aggressive frequency reuse, dense orbital constellations, and adjacent-channel leakage. Tunable Spectral Selectivity: Laser-controlled energy level coupling, Rydberg state-dependent dipole transitions, and sub-kHz spectral resolution.
Hardware Instability: Thermal drift under vacuum conditions and oscillator drift. Quantum Characteristics: Quantum-coherent field transduction and atomic energy levels with thermal insensitivity.
@@ II-A2 Representative Implementations
To realize the above-mentioned quantum sensing processes, a typical solution is the two-photon excitation scheme, as demonstrated in Fig. 1(b). An ensemble of Rydberg atoms in the form of the atomic vapor is filled up with a glass cell. A pair of laser beams, termed as ‘probe’ and ‘coupling’, counter-propagate through the vapor cell and form a spatially overlapped detection region. Under the joint excitation of the probe and coupling beams, the atoms of this region are excited to the Rydberg state, realizing the quantum state preparation. During this atom-light interaction process, a special but useful physical phenomenon, known as the Rydberg EIT, is generated, which renders a transparent window of the atomic vapor over a narrow range of frequencies of the absorption spectrum. Specifically, by overlapping the probe and coupling beams, the probe beam that should have been absorbed by the atomic vapor becomes transmissible as if the atomic vapor is transparent, as depicted in Fig. 1(d)(ii).
@@ II-A4 Mathematical Theories
RAQRs belong to a true interdisciplinary field, involving quantum sensing, atom-light interaction, photodetection, RF, and baseband signal processing. To completely characterize the holistic system, an integration of distinctive mathematical theories is required. Specifically, the Lindblad master equation is typically employed to describe the dynamic evolution of the quantum eigenstates and coherence between different eigenstates, providing a microscopic characterization of the quantum sensing mechanism. The microscopic coherence can be further associated with the macroscopic susceptibility of the atomic vapor that directly affects the laser beam propagation. Therefore, the atom-light interaction can be characterized by exploiting Maxwell’s wave equation in a medium, i.e., in the atomic vapor. Furthermore, the characterization of the photodetection relies on the theory of the photoelectric effect. By integrating all these theories into a communication and sensing framework, an equivalent baseband signal model of the end-to-end transceiver was constructed in [7], paving the way for applying diverse signal processing algorithms.
@@ II-B Natural Benefits of RAQRs for Satellite Applications
Based on the characteristics mentioned above, the integration of RAQRs with satellite systems offers a compelling solution to key satellite communication challenges, primarily by leveraging four natural matches.
High Sensitivity: Compared to traditional RF receivers, RAQRs demonstrate ultra-high sensitivity in detecting RF fields. Their sensitivity can reach the order of $\text{nv}\textfractionsolidus\text{cm}/\sqrt{\text{Hz}}$ [11], surpassing the typical sensitivity limit of conventional metal antennas. This exceptional sensitivity implies that RAQR-assisted satellites are capable of directly receiving extremely weak signals from ground users without large antenna apertures or high-gain amplification.
Miniaturized Architecture: Unlike conventional antennas, the effective size of RAQRs is independent of the signal wavelength, as RAQRs receive RF signals through Rydberg atoms excited by laser beams within a vapor cell. Compared to conventional antennas (300mm at 1 GHz), the effective region of RAQRs is between 0.1mm and 10 mm [12], enabling extreme miniaturization. As a result, RAQRs offer a compact and lightweight alternative to classical RF front-ends.
@@
Tunable Spectral Selectivity: QAQRs offer a highly advantageous blend of uniquely tunable bands and sharp spectral selectivity [13], making them exceptionally suitable for satellite communication applications. By adjusting laser frequencies and engaging specific Rydberg energy levels, these receivers achieve quasi-continuous coverage spanning MHz to THz frequencies with instantaneous sub-kilohertz spectral resolution. This sharp selectivity enables precise differentiation of adjacent-band signals, reducing adjacent-band interference and enhancing spectrum efficiency accordingly.
Environmental Benefits: The ultra-high vacuum environment of space, such as within satellite payload chambers, unlocks the full potential of RAQRs by significantly reducing atomic decoherence. In these conditions, gas collisions are significantly reduced, extending Rydberg state lifetimes by orders of magnitude [14]. Moreover, suppressed black-body radiation under deep vacuum and cryogenic conditions minimizes thermal excitations and spontaneous decay [15], enhancing spectral purity and coherence.
These benefits collectively enable ultra-narrow linewidths, quantum-limited sensitivity, and high-precision, low-noise communication for satellite systems, thereby enhancing the reliability of ground-satellite links.
@@ III-A Hybrid Transceiver Architecture
Figure 2 shows the elaborately designed hybrid architecture for the RAQRs and RF transceiver. Particularly, the devised transceiver can be divided into two parts, including a conventional RF transmitter and a hybrid receiver. A conventional RF transmitter typically consists of a baseband processing unit, a modulator, and a power amplifier. A hybrid receiver consists of two main components: an RF antenna receiver and a RAQR. The RF antenna receiver comprises a low-noise amplifier, a demodulator, and a signal processor, while the RAQR consists of a Rydberg atomic quantum sensor, a photodetection unit, a down-conversion unit, and a signal processor.
The equivalent baseband signal is further detailed in [10]. Using the received signals from RAQR and the RF receiver simultaneously, the following two models can be operated, as illustrated in Fig. 3.
@@
Wideband Reception/Transmission: Although RAQRs are inherently limited to narrowband reception, they can provide highly accurate channel state information (CSI) within selected frequency bands. By applying interpolation techniques, the partial CSI can be extended across the entire signal bandwidth, thereby improving the performance of conventional RF transceivers in both detection and transmission via hybrid processing. Based on the quantum parameters of Table I in [7], and assuming a propagation distance of 500 km with $M=100$ Rydberg sensors, we apply the minimum mean square error method to evaluate the mean square error of channel estimation and the bit error rate (BER) performance under QPSK modulation, as illustrated in Fig. 3(b). Owing to the intrinsic high sensitivity of Rydberg atoms, the accuracy of channel estimation is significantly improved, leading to over 20 dB enhancement in wideband signal detection performance. This result confirms that direct ground-to-satellite links can be substantially enhanced.
@@ III-B Scalable RAQR-Based Satellite Networks
To enhance the spatial diversity and improve the coverage, we further investigate the scalable RAQR-based satellite networks as shown in Fig. 4(a), where multiple satellites and ground terminals are equipped with RAQRs operating collaboratively within a distributed architecture. Specifically, each satellite can perform local detection or transmit the received signal to the central processing satellite (head of the dynamic cluster) for global detection.
@@
Based on the above-mentioned parameters, Figure 4(b) depicts the BER performance with two types of detection. As expected, global detection achieves near-optimal performance by effectively suppressing inter-satellite interference. By contrast, as the number of satellites increases, the BER of local detection tends to plateau. This performance saturation arises from the limited capability of localized processing to mitigate interference originating from distant satellites, thereby leading to a performance bottleneck. Notably, compared to conventional RF-enabled satellite systems, RAQR-based architectures offer enhanced direct ground-to-satellite access and represent a promising solution for next-generation satellite communication networks.
@@ IV Performance Evaluations of Direct Access
Next, we introduce the RAQRs into classical satellite networks and reveal the potential of RAQR-enabled satellite networks by detailing the significant improvement on the achievable rate, coverage area, and sensing accuracy.
A four-level electron transition of 6S 1/2 →6P 3/2 →47D 5/2 →48P 3/2 is considered in our simulations, corresponding to the detection of RF signals having a carrier frequency of $f_{c}=6.9458$ GHz. The parameters of electron transitions, laser beams, RF signals, and electronic components follow the same values as those listed in TABLE I of [7]. It is emphasized that our configurations are consistent with the physics experiments and 3GPP specification. We employ $100$ Rydberg sensors equipped on the satellite to form a uniform linear array for receiving the signal transmitted from a single-antenna ground station. For simplicity, we assume a line-of-sight propagation for the ground-to-satellite transmission. The FSPL is given by $20\log_{10}\left(c/(4\pi)\right)+20\log_{10}\left(1/d\right)+20\log_{10}\left(1/f_{c}\right)$ in dB, where $c$ represents the speed of light and $d$ is the distance.
@@ IV-B Coverage Area
As illustrated in Fig. 5(b), owing to the ultra-high sensitivity, RAQRs enable satellite-based reception of ground-transmitted signals over distances as long as $2000$ km. By comparison, conventional RF antennas operating under the same signal strength constraints (10 dB) are limited to ranges of around $100$ km. This substantial extension (around 1900 km) in communication range demonstrates the potential of RAQRs to significantly enhance the reliability of the direct satellite-to-ground link, providing a new paradigm for future satellite communication systems.

### v2

@@ Abstract
Ground-satellite links for 6G networks face critical challenges, including severe path loss, tight size‑weight‑power limits, and congested spectrum, all of which significantly hinder the performance of traditional radio frequency (RF) front ends. This article introduces the Rydberg Atomic Quantum Receiver (RAQR) for onboard satellite systems, a millimetre‑scale front end that converts radio fields to optical signals through atomic electromagnetically induced transparency. RAQR’s high sensitivity and high frequency selectivity have the potential to address link‑budget, payload, and interference challenges while fitting within space constraints. Theoretically, a hybrid atomic–electronic design that is supported by a consistent signal model achieves spectral efficiency exceeding 6 bit/s/Hz, extends coverage by up to 1000 km, and improves sensing accuracy by two orders of magnitude relative to conventional RF receivers. The paper concludes with integration strategies, distributed‑satellite concepts, and open research challenges for bringing RAQR‑enabled satellite payloads into service.
@@ I Introduction
Spectrum Congestion and Interference Vulnerability: Satellite frequency bands are heavily shared among civil, commercial, and governmental users (e.g., L, S, C, Ku bands). To increase throughput, aggressive frequency reuse strategies are common. These strategies, in turn, introduce substantial co-channel and adjacent-channel interference from other ground stations or satellite systems [5]. Moving to higher frequency bands (e.g., Ka or Q/V bands) offers more bandwidth but brings new design challenges for onboard receivers. These challenges include increased rain fade, stricter hardware linearity requirements, and elevated thermal instability in front-end components. Furthermore, ground-based jamming poses a direct threat to uplink integrity, which requires onboard receivers with robust interference rejection.
Overcoming these intertwined constraints calls for a receive architecture on the satellite that combines very high sensitivity, compact physical dimensions, and minimal power consumption while maintaining interference tolerance. This motivates the investigation of radically different front-end architectures capable of operating under extreme power, size, and spectral isolation constraints, particularly for the reception of ground-originated signals. Quantum-enhanced sensing technologies represent a particularly promising direction. One such approach is the Rydberg atomic quantum receiver (RAQR), which utilizes highly excited Rydberg states of atoms to transduce RF fields from the ground segment into optical signals [6]. These receivers can directly demodulate RF signals via electromagnetically induced transparency (EIT) without requiring conventional components such as local oscillators or RF mixers. Their unique properties, such as wide frequency tunability, ultra-high field sensitivity to weak signals, inherent optical isolation from thermal noise, and the potential for chip-scale integration, make RAQRs an intriguing candidate for next-generation spaceborne front-ends [6].
@@
TABLE I: RAQR Capabilities Relevant to Satellite Challenges
Satellite Communication Challenges RAQR-Based benefits Practical challenges and potential solutions
Link Budget Constraints: Severe FSPL (150 - 200 dB), constrained power, and limited antenna gain. High Sensitivity: Ultra-high dipole moment between Rydberg transitions and limited thermal noise impacts [6, 7]. Challenges: Experimental sensitivity related to atomic density, temperature, detection, and laser drift. Solutions: Frequency-locked lasers, ovenized cell, and thermal management.
Payload SWaP Constraints: Constrained volume and mass budgets, limited solar panel area, and power generation. Miniaturized Architecture: Flexible extension for multi-functional receivers using a single vapor cell, wavelength-independent sensing volume, and no traditional RF frontend [6]. Challenges: Stabilized laser provisioning, micro vapor-cell packaging and alignment, and frequency calibration. Solutions: Centralized stabilized master laser, module multiplexing, and micro-integrated waveguide .
Spectrum Congestion and Interference: Aggressive frequency reuse, dense orbital constellations, and adjacent-channel leakage. Tunable Spectral Selectivity: Laser-controlled energy level coupling, Rydberg state-dependent dipole transitions, and sub-kHz spectral resolution [8]. Challenges: Limited instantaneous bandwidth and discrete frequency coverage. Solutions: Multi-band or multi-carrier transmission, multi-species/multi-series state, and atomic heterodyne mode.
Hardware Instability: Thermal drift under vacuum conditions and oscillator drift. Quantum Characteristics: Quantum-coherent field transduction and atomic energy levels with thermal insensitivity Challenges: Temperature, frequency drift, blackbody radiation-induced dephasing, and power broadening. Solutions: Multi-zone ovenized cell, automatic drift compensation, and periodic recalibration.
@@ II-A2 Representative Implementations
Fig. 1(b) illustrates the above-mentioned quantum sensing processes of the two-photon excitation scheme. An ensemble of Rydberg atoms in the form of the atomic vapor is filled up with a glass cell. A pair of laser beams, termed as ‘probe’ and ‘coupling’, counter-propagate through the vapor cell and form a spatially overlapped detection region. Under the joint excitation of the probe and coupling beams, the atoms of this region are excited to the Rydberg state, realizing the quantum state preparation. During this atom-light interaction process, a special but useful physical phenomenon, known as the Rydberg EIT, is generated, which renders a transparent window of the atomic vapor over a narrow range of frequencies of the absorption spectrum. Specifically, by overlapping the probe and coupling beams, the probe beam that should have been absorbed by the atomic vapor becomes transmissible as if the atomic vapor is transparent, as depicted in Fig. 1(d)(ii).
@@ II-A4 Mathematical Theories
RAQRs lie at the intersection of quantum sensing, atom–light interaction, photodetection, and classical RF/baseband signal processing, requiring a unified modelling framework that bridges microscopic quantum physics with macroscopic signal observables. The modelling can be understood in a stepwise manner. Starting from the Beer–Lambert law, one can describe how the probe laser experiences absorption and phase shift when propagating through an atomic vapor. These absorption and dispersion effects are governed by the medium’s electric susceptibility, which in turn is derived from the Maxwell–Bloch equations linking Maxwell’s electromagnetic wave equation with the atomic density matrix. The atomic density matrix itself evolves according to the Lindblad master equation, which captures the coherent and dissipative dynamics among quantum states. The resulting optical response modulates the probe beam intensity, which is then converted into an electrical signal via the photoelectric effect in the photodetector. Integrating these layers—from quantum coherence to optical response and finally to baseband signal—enables the construction of an equivalent end-to-end signal model for the RAQR transceiver [6], providing a clear pathway to analyze system performance and apply advanced signal processing algorithms.
@@ II-B Natural Benefits of RAQRs for Satellite Applications
Although integrating the RAQR into a satellite platform remains practically challenging, this approach offers compelling advantages for addressing key satellite communication challenges, primarily through four intrinsic compatibilities, as summarized in Table I.
High sensitivity: Due to practical limitations of the experimental setup, RAQRs did not demonstrate comprehensive superiority over conventional receivers. As revealed in [11], the sensitivity highly depends on the experimental configuration, frequency, the detection schemes, and the suppression of practical impairments. In particular, resonant operation enhances the field sensitivity from 23 to 1.59 $\mathrm{\mu V/cm/\sqrt{Hz}}$ at 10.22 GHz and room temperature [12]. Recently, the $\mathrm{nV/cm/\sqrt{Hz}}$ field-equivalent sensitivities at 8.57 GHz can be achieved by using RF heterodyne detection [13]. Promisingly, the recent theoretical analyses in [6, 7] and experimental study [14] have exhibited strong evidence for supporting the superiority of RAQRs. This exceptional sensitivity implies that RAQR-assisted satellites are capable of directly receiving extremely weak signals from ground users without large antenna apertures or high-gain amplification.
Miniaturized Architecture: Unlike conventional antennas, the effective size of RAQRs is independent of the signal wavelength, as RAQRs receive RF signals through Rydberg atoms excited by laser beams within a vapor cell [6]. Existing research has shown that it is possible to replace an RF antenna with a Rydberg atomic sensor. Although components such as laser arrays and frequency-stabilization hardware add mass and volume, these SWaP costs can be reduced through chip-scale photonic integration. Additionally, laser and signal-processing modules can be multiplexed with the spacecraft’s laser communications terminal and can share its stabilized optical platform, power, and thermal management services, thereby mitigating SWaP constraints. As a result, RAQRs have the potential to provide a compact alternative, although meeting SWaP targets is still challenging, especially for the laser subsystem.
@@
Tunable Spectral Selectivity: RAQRs offer a highly advantageous blend of uniquely tunable bands and sharp spectral selectivity [8], making them exceptionally suitable for satellite communication applications. By adjusting laser frequencies and engaging specific Rydberg energy levels, these receivers achieve quasi-continuous coverage spanning MHz to THz frequencies with instantaneous sub-kilohertz spectral resolution. This sharp selectivity enables precise differentiation of adjacent-band signals, reducing adjacent-band interference and enhancing spectrum efficiency accordingly.
Vacuum Benefits: Even though day-night cycling leads to an unstable temperature for Rydberg atoms, the satellite vacuum environment facilitates stable RAQR operation [15]. Particularly, the natural vacuum space enables easier thermal control, reduces pressure and acoustic perturbations, and lowers gas-collision rates in the vapor cell, which enhances the optical response. Moreover, with appropriate thermal shielding, the interference posed by unexpected radiation can be mitigated, leading to higher sensitivity. In other words, the natural vacuum environment provides a favorable environment that can reduce some perturbations, though practical challenges such as thermal cycling, radiation, and laser stability remain significant.
Although integration still faces outstanding engineering challenges, including hardware implementation, thermal management, and laser stabilization, these characteristics have the potential to support narrow linewidths and high sensitivity, though achieving this in a space-qualified system remains an open engineering challenge.
@@ III-A Hybrid Transceiver Architecture
Fig. 2 shows the elaborately designed hybrid architecture for the RAQRs and RF transceiver. Particularly, the devised transceiver can be divided into two parts, including a conventional RF transmitter and a hybrid receiver. A conventional RF transmitter typically consists of a baseband processing unit, a modulator, and a power amplifier. A hybrid receiver consists of two main components: an RF antenna receiver and a RAQR. The RF antenna receiver comprises a low-noise amplifier, a demodulator, and a signal processor, while the RAQR consists of a Rydberg atomic quantum sensor, a photodetection unit, a down-conversion unit, and a signal processor.
The equivalent baseband signal is further detailed in [10]. Based on the received signals from RAQR and the RF receiver simultaneously, Fig. 3(a)(i) and Fig. 3(a)(ii) illustrate wideband reception and narrowband reception, respectively.
@@
Wideband Reception/Transmission: Although RAQRs are inherently limited to narrowband reception, they can provide highly accurate channel state information (CSI) within selected frequency bands. By applying interpolation techniques, the partial CSI can be extended across the entire signal bandwidth, thereby improving the performance of conventional RF transceivers in both detection and transmission via hybrid processing. Using the quantum parameters of Table I in [6] for the four-level electron transition of 6S 1/2 →6P 3/2 →47D 5/2 →48P 3/2 , and assuming that the noise comes from three parts, including thermal noise, quantum projection noise, and photon shot noise, we evaluate the channel estimation performance by using the minimum mean square error method and then investigate the bit error rate (BER) performance under QPSK modulation, as illustrated in Fig. 3(b). Owing to the intrinsic high sensitivity of Rydberg atoms, the accuracy of channel estimation is significantly improved, leading to over 20 dB projected enhancement under the idealized assumptions. From a theoretical perspective, this result indicates that RAQRs have the potential to address the direct ground-to-satellite connectivity challenge.
@@ III-B Scalable RAQR-Based Satellite Networks
To enhance the spatial diversity and improve the coverage, we further investigate the scalable RAQR-based satellite networks. As shown in Fig. 4(a), multiple satellites and ground terminals are equipped with RAQRs operating collaboratively within a distributed architecture. Specifically, each satellite can perform local detection or transmit the received signal to the central processing satellite (head of the dynamic cluster) for global detection.
@@
Using the above-mentioned parameters, Fig. 4(b) provides a theoretical projection of the BER performance for the two detection schemes. As expected, global detection achieves near-optimal performance by effectively suppressing inter-satellite interference. By contrast, as the number of satellites increases, the BER of local detection tends to plateau. This performance saturation arises from the limited capability of localized processing to mitigate interference originating from distant satellites, thereby leading to a performance bottleneck. Notably, compared to conventional RF-enabled satellite systems, RAQR-based architectures offer enhanced direct ground-to-satellite access and represent a promising solution for next-generation satellite communication networks.
@@ IV Performance Evaluations of Direct Access
Next, we present a theoretical projection of integrating RAQRs into classical satellite networks to explore their potential benefits, including possible enhancements in achievable rate, coverage area, and sensing accuracy.
A four-level electron transition of 6S 1/2 →6P 3/2 →47D 5/2 →48P 3/2 is considered in our simulations, corresponding to the detection of RF signals having a carrier frequency of $f_{c}=6.9458$ GHz. The parameters of electron transitions, laser beams, RF signals, and electronic components follow the same values as those listed in TABLE I of [6]. These configurations are consistent with the physics experiments and 3GPP specification. Based on these parameter settings, by assuming that there are 100 Rydberg sensors equipped on the satellite, we investigate the benefits of the RAQR-based satellite over a conventional one. Furthermore, we assume a line-of-sight propagation for the ground-to-satellite transmission. The FSPL is given by $20\log_{10}\left(c/(4\pi)\right)+20\log_{10}\left(1/d\right)+20\log_{10}\left(1/f_{c}\right)$ in dB, where $c$ represents the speed of light and $d$ is the distance.
@@ IV-B Coverage Area
As illustrated in Fig. 5(b), owing to the ultra-high sensitivity, RAQRs enable satellite-based reception of ground-transmitted signals over distances as long as $2000$ km. By comparison, conventional RF antennas operating under the same signal strength constraints (10 dB) are limited to ranges of around $100$ km. Furthermore, for achieving the same received SNR, the energy efficiency relying on RAQR is higher than that based on RF receiver. This substantial extension (around 1900 km) and high efficiency in communication demonstrates the potential of RAQRs to significantly enhance the reliability of the direct satellite-to-ground link, providing a new paradigm for future satellite communication systems.